OpenStudy (anonymous):

solve by factoring 12x^2+9=24x

OpenStudy (anonymous):

12x^2-24x+9=0 which can be reduced as 4x^2-8x+3=0 4x^2-6x-2x+3=0 2x(2x-3)-1(2x-3)=0 (2x-1)(2x-3)=0 so x=1/2 and x=3/2